carriefreeman:

Alan deposited $300 in an account that pays 6% interest compounded continuously. Approximately how long will it take for Alan's money to triple?

jhonyy9:

how much is this 6% of $300 ? can you calcule it ? how much will be the Alan's money tripled ?

Mercury:

continuous compounding formula: A = Pe^(rt) where P is the principal, e is a constant (approx: 2.71828), r is the interest rate as a decimal (0.06) and t is time (which you'll solve for) since it wants to know how long it'll take for the $ to triple, A = 3P substituting that into the equation, we get: 3P = Pe^(0.06t) notice how P cancels out on both sides: 3 = e^(0.06t) solve for t.